### v1

@@ 3.2 Implementation Details
SwinUnetr [15] 44.408/0.976 44.979/0.979 45.417/0.981 46.059/0.983 46.774/0.986
MambaMIR [12] 44.028/0.968 44.473/0.973 45.079/0.976 45.487/0.978 45.914/0.980
UNN [21] 44.504/0.972 45.185/0.977 45.663/0.980 46.078/0.984 46.646/0.984
HF-ResDiff [18] 44.728/0.979 45.243/0.982 45.870/0.984 46.174/0.986 47.004/0.984
@@
CNN [17] 39.570/0.967 39.532/0.961 40.049/0.963 40.276/0.966 40.680/0.968
SwinUnetr [15] 44.119/0.966 44.560/0.969 44.866/0.971 45.171/0.975 45.783/0.979
MambaMIR [12] 44.028/0.968 44.429/0.970 44.808/0.970 45.207/0.974 45.844/0.979
UNN [21] 44.759/0.969 45.219/0.971 45.432/0.975 45.765/0.977 46.384/0.981
HF-ResDiff [18] 44.219/0.964 44.632/0.968 45.217/0.971 45.493/0.974 46.287/0.979

### v2

@@ 3.2 Implementation Details
SwinUnetr [15] 44.408/0.976 44.979/0.979 45.417/0.981 46.059/0.983 47.115/0.986
MambaMIR [12] 44.349/0.978 45.801/0.982 45.830/0.983 46.043/0.985 46.774/0.987
UNN [21] 44.504/0.972 45.185/0.977 45.663/0.980 46.071/0.982 46.646/0.984
HF-ResDiff [18] 44.728/0.979 45.243/0.982 45.677/0.984 46.174/0.986 47.004/0.984
@@
CNN [17] 39.570/0.967 39.832/0.961 40.049/0.963 40.276/0.965 40.680/0.968
SwinUnetr [15] 44.119/0.966 44.560/0.969 44.866/0.971 45.171/0.973 45.783/0.976
MambaMIR [12] 44.028/0.968 44.429/0.970 44.810/0.972 45.207/0.974 45.844/0.977
UNN [21] 44.759/0.969 45.019/0.973 45.432/0.975 45.765/0.977 46.496/0.980
HF-ResDiff [18] 44.219/0.964 44.632/0.968 45.081/0.971 45.493/0.974 46.287/0.977
@@ 3.3 Comparison with State-of-the-Art Methods
Figure 4: Quantitative comparison with state-of-the-art methods at 2% dose level.